### v2

@@ 3.2.1 Integration into Our Framework
Traditional self-play algorithms can be incorporated into our proposed framework (Algo. 1) with the following settings. First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy. Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initialized randomly, learning from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$. Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation. Only the MSS for prioritized fictitious self-play (described in Sec. 3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes such as NE.
@@
In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,\leavevmode\nobreak\ \forall i$.
@@
The policy population gradually increases over time. Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are inactive policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
@@ 3.2.2 Vanilla Self-play
In vanilla self-play [108], agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (5)
@@ 3.2.3 Fictitious Self-play
Fictitious Play (FP) [11] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-Play (FSP) [50] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-Play (NFSP) [51] is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
In more recent approaches [73], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (6)
@@ 3.2.4 $\delta$-uniform Self-play
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (7)
@@ 3.2.5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-Play (PFSP)  [129] leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$. The MSS of PFSP is given by Algo. 5. The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics. A similar MSS in  [7] can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
Algorithm 5 the PFSP Meta-strategy Solver.
@@
2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i$
@@ 3.3.1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [73] in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria [92, 96, 86]. Lastly, we also simply follow Assumption 1. We can derive the Corollary 2 using a similar proof as Corollary 1.
@@ 3.3.2 Double Oracle
Double Oracle (DO) [92] can be only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 the NE-based Meta-strategy Solver.
@@ 3.3.3 PSRO
PSRO [73] extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
@@ 3.3.4 PSRO Variants
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. By establishing a hierarchical structure of RL workers, Pipeline PSRO [87] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO [151] introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO. Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) [89] allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games. Online Double Oracle (ODO) [35] integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff. Anytime PSRO [90] and Self-play PSRO [88] are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO [97] has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games [32, 31] and the selection problem in NE [44], Müller et al. put forward $\alpha$-PSRO [96]. Instead of NE, $\alpha$-rank [101], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, [86] proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games. It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
@@
Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful. [4] introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. [103] uses determinantal point process [69] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO. Similarly,  [83] introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO [142] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
@@ 3.3.5 R-NaD
Although R-NaD [105] is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo. 3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics [124] until convergence to a fixed point. The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (5). This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
@@ 3.4 Ongoing-training-based Series
In the PSRO series of algorithms, two key challenges arise. First, truncating the ABR operators during each iteration is often necessary when operating with a limited budget, introducing sub-optimally trained responses into the population. Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents [82]. To address these challenges, the ongoing-training-based series of algorithms emphasizes the repeated ongoing training of all policies. In other words, instead of only inactive policies being selected for training, all active policies are also likely to be chosen for training.
@@ 3.4.2 FTW
Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent [60] is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:

### v3

@@ 3.2.1 Integration into Our Framework
Traditional self-play algorithms can be incorporated into our proposed framework (Algo. 1) with the following settings. First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy. Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initialized randomly, learning from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$. Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation. Only the MSS for prioritized fictitious self-play (described in Sec. 3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes like NE.
@@
In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,~{}\forall i$.
@@
The policy population gradually increases over time. When policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are inactive policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
@@ 3.2.2 Vanilla Self-play
In vanilla self-play (Samuel, 2000), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,~{}\text{if}~{}j=i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (5)
@@ 3.2.3 Fictitious Self-play
Fictitious Play (FP) (Brown, 1951) is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-Play (FSP) (Heinrich et al., 2015) is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-Play (NFSP) (Heinrich & Silver, 2016) is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
In more recent approaches (Lanctot et al., 2017), random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},~{}\text{if}~{}j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (6)
@@ 3.2.4 $\delta$-uniform Self-play
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},~{}\text{if}~{}\lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right.,$ (7)
@@ 3.2.5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-Play (PFSP)  (Vinyals et al., 2019) leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}~{}\text{beats}~{}\pi_{j})$. The MSS of PFSP is given by Algo. 5. The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics. A similar MSS in  Berner et al. (2019) can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
Algorithm 5 PFSP meta-strategy solver.
@@
2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},~{}\text{if}~{}j\leq i$
@@ 3.3.1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021). Lastly, we also simply follow Assumption 1. We can derive the Corollary 2 using a similar proof as Corollary 1.
@@ 3.3.2 Double Oracle
Double Oracle (DO) (McMahan et al., 2003) can be only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 NE-based meta-strategy solver.
@@ 3.3.3 PSRO
PSRO (Lanctot et al., 2017) extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
@@ 3.3.4 PSRO Variants
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. By establishing a hierarchical structure of RL workers, Pipeline PSRO (McAleer et al., 2020) achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO (Zhou et al., 2022) introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO. Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) (McAleer et al., 2021) allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games. Online Double Oracle (ODO) (Dinh et al., 2022) integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff. Anytime PSRO (McAleer et al., 2022b) and Self-play PSRO (McAleer et al., 2022a) are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO (Muller et al., 2022) has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013), Müller et al. put forward $\alpha$-PSRO (Muller et al., 2020). Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, Marris et al. (2021) proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games. It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
@@
Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful. Balduzzi et al. (2019) introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. Perez-Nieves et al. (2021) uses determinantal point process (Kulesza et al., 2012) to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO. Similarly,  Liu et al. (2021) introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO (Yao et al., 2023) defines a diversity metric named population exploitability that helps to achieve a full-game NE.
@@ 3.3.5 R-NaD
Although R-NaD (Perolat et al., 2021) is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo. 3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor & Jonker, 1978) until convergence to a fixed point. The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (5). This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
@@ 3.4 Ongoing-training-based Series
In the PSRO series of algorithms, two key challenges arise. First, truncating the ABR operators during each iteration is often necessary when operating with a limited budget, introducing sub-optimally trained responses into the population. Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents (Liu et al., 2022c). To address these challenges, the ongoing-training-based series of algorithms emphasizes the repeated ongoing training of all policies. In other words, instead of only inactive policies being selected for training, all active policies are also likely to be chosen for training.
@@ 3.4.2 FTW
Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent (Jaderberg et al., 2019) is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS: